OpenStudy (anonymous):

Which best explains why the equation 3x+8=3x-5 has no solutions A. The equation has a variable on both sides. B.the equation is not true for any value of x C.the equation is true for any value of x

jimthompson5910 (jim_thompson5910):

you can think of 3x as z ie, z = 3x

jimthompson5910 (jim_thompson5910):

3x+8=3x-5 would then turn into z+8 = z-5

OpenStudy (anonymous):